Jared borrowed $20,000 with a promise to repay the loan in 6 years with a uniform monthly payment and a single payment of $2,000 at the end of six years at a nominal interest rate of 12% per year.A) What is the amount of each payment? B) What is the amount of interest paid in the first payment? C) What will be the loan balance immediately after the 48th payment? D) What is semi annually effective interest rate?
Answer #1

1.
=-PMT(12%/12,6*12,20000,-2000)=$371.90

2.
=20000*12%/12=200

3.
=-FV(12%/12,48,PMT(12%/12,6*12,20000,-2000),20000)=9475.833725

4.
=(1+12%/2)^2-1=12.36%